OpenStudy (anonymous):

did i work this our correctly? solve; r=2 thirds*t+v, for t. -v -v r-v=two thirds*t (multiplied be three) *3 *3 r-v*3=2t (divided be two) r-v*3(over a fraction bar) 2 (under the fration bar) =t PLZ HELP! I need to get a 100 on this!

OpenStudy (anonymous):

solve; \[r=\frac{ 2 }{ 3 }t+v\], for t.

OpenStudy (anonymous):

r = 2 t/3 + v 2 t/3 = r - v => t = (3/2 ) ( r -v)

OpenStudy (anonymous):

i got \[t=\frac{ r-v*3 }{ 2 }\] when i was compleatly done.

OpenStudy (anonymous):

You need the bracket: t = 3 ( r -v) /2

OpenStudy (anonymous):

she never taught us that, tho..

OpenStudy (anonymous):

the way i typed it is how we end........(me confused)

OpenStudy (anonymous):

@Chlorophyll , does it mean the same thing??? what u put and what i put?

OpenStudy (anonymous):

@TuringTest @Callisto do u think u can help? did i work this correctly?

OpenStudy (anonymous):

NO, they're different!

OpenStudy (anonymous):

OMG! How?!

OpenStudy (anonymous):

t = 3 ( r -v) /2 = ( 3r - 3v ) /2 Your answer: t = ( r - 3v} /2

OpenStudy (anonymous):

why multiply v and r by three?????! i dont get it!

OpenStudy (anonymous):

r = 2 t/3 + v 2 t/3 = r - v Multiply both sides for 3/2 :

OpenStudy (anonymous):

Your post " r=2 thirds*t+v " r = 2 t/3 + v In case: r = 2 thirds ( t + v), then the result is different!

OpenStudy (anonymous):

 dont mind the rectangle or the line over to the side. i was messeing around a bit..